### v1

# SyntheT2C: Generating Synthetic Data for Fine-Tuning Large Language Models on the Text2Cypher Task

Abstract
Integrating Large Language Models (LLMs) with existing Knowledge Graph (KG) databases presents a promising avenue for enhancing LLMs’ efficacy and mitigating their “hallucinations”. Given that most KGs reside in graph databases accessible solely through specialized query languages (e.g., Cypher), there exists a critical need to bridge the divide between LLMs and KG databases by automating the translation of natural language into Cypher queries (commonly termed the “Text2Cypher” task). Prior efforts tried to bolster LLMs’ proficiency in Cypher generation through Supervised Fine-Tuning. However, these explorations are hindered by the lack of annotated datasets of Query-Cypher pairs, resulting from the labor-intensive and domain-specific nature of annotating such datasets. In this study, we propose SyntheT2C, a methodology for constructing a synthetic Query-Cypher pair dataset, comprising two distinct pipelines: (1) LLM-based prompting and (2) template-filling. SyntheT2C facilitates the generation of extensive Query-Cypher pairs with values sampled from an underlying Neo4j graph database. Subsequently, SyntheT2C is applied to two medical databases, culminating in the creation of a synthetic dataset, MedT2C. Comprehensive experiments demonstrate that the MedT2C dataset effectively enhances the performance of backbone LLMs on the Text2Cypher task. Both the SyntheT2C codebase and the MedT2C dataset will be released soon.
SyntheT2C: Generating Synthetic Data for Fine-Tuning Large Language Models on the Text2Cypher Task
1 Introduction
Figure 1: SyntheT2C builds synthetic data with two pipelines to SFT LLMs so that their performance on Text2Cypher task is enhanced.
Knowledge Graphs (KGs) constitute vital reservoirs of information within the Retrieval-Augmented Generation (RAG) paradigm Lewis et al. (2020) of Large Language Models (LLMs). Distinguished from other information sources, KGs boast structured and meticulously curated data, rendering them conducive to seamless updates and rectifications. Such attributes position KGs as pivotal instruments for mitigating issues of knowledge cutoff and “hallucinations” within LLMs. Notably, KGs have long served as a core in numerous knowledge-intensive products and applications Kertkeidkachorn et al. (2023); Cui et al. (2024); Xu et al. (2020). With the advent of LLMs, many researchers have focused on synergizing KGs with LLMs following the RAG framework. The inherent fidelity and adaptability of KGs make them practical assets for deployment in production environments, and also catapult KGs to the forefront of academic research.
While KGs represent invaluable repositories of reference information, their efficient utilization remains a formidable challenge. Early methodologies involved direct extraction of triplets from KGs, subsequently integrating these text-form triplets directly into the prompts of LLMs Fatemi et al. (2023). However, this approach often fails to concurrently preserve both semantic and structural nuances inherent within the KG. An alternative approach involves querying existing graph databases just like human users, promising accurate and interpretable results. Nonetheless, the primary impediment lies in the LLM’s ability to formulate correct and executable queries. To address this limitation, numerous query generation tools or methodologies Zhang et al. (2022); Abdelaziz et al. (2021); Shen et al. (2023) are proposed, aiming to translate human users’ natural language queries into query languages. This task assumes paramount importance for LLM development for two pivotal reasons: (1) it empowers LLMs to consistently produce reliable queries, thereby augmenting their utilization of existing KG databases to address knowledge deficits; (2) it facilitates human interaction with KG databases through natural language, substantially lowering the barrier to entry for KG database utilization. Among the spectrum of query generation research, the sub-task of translating natural language into the Cypher Francis et al. (2018) query language for Neo4j Neo4j (2012) databases stands out as a prominent research focus. This prominence is attributed to two key factors. Firstly, Neo4j is a widely adopted solution for constructing KG databases, positioning Cypher as an essential tool for accessing these extensive repositories. Secondly, Cypher is a query language specifically designed for querying graph structures, offering significantly faster performance than other query languages, such as SQL, when processing graph data. Consequently, our work centers on this sub-task, commonly termed as “Text2Cypher” (T2C).
A similar task to the Text2Cypher task is the “Text2SQL” task, wherein researchers endeavor to translate natural language sentences into SQL queries. Leveraging manually annotated datasets like SPIDER Yu et al. (2019), numerous methodologies have emerged, including SpCQL Guo et al. (2022) and SQLNet Xu et al. (2017). Conversely, scant attention has been directed towards the Text2Cypher task. Existing approaches typically resort to decomposing the original query into smaller components and translating each part separately. For instance, $R^{3}$-NL2SQL Zhou et al. (2023b) partitions the query generation process into CRUD keywords prediction, clause selection, and object type identification. Despite the success of these methods, adapting them to a specific KG database demands substantial extra effort. With the rise of LLMs, using LLMs for Cypher query generation appears promising. Notably, to the best of our knowledge, no endeavors have explored the potential application of LLMs to the Text2Cypher task. Our work aims to bridge this gap in the literature.
The Cypher writing performance of vanilla LLMs is not satisfactory. To improve it, we employ SFT, which necessitates a dataset of Question-Query pairs. However, creating such a dataset is challenging as it requires both domain-specific knowledge of the KG’s content and expertise in Cypher’s syntax. Consequently, there is currently no annotated dataset for the Text2Cypher task. To overcome this obstacle, we introduce SyntheT2C, a method designed to produce high-quality synthetic Question-Cypher pairs through two distinct pipelines: LLM-based prompting and template-filling (as shown in Figure 1). The LLM-based prompting pipeline aims to generate Cypher queries with greater semantic flexibility, while the template-filling pipeline focuses on producing syntactically complex Cypher queries. The generated Question-Query pairs undergo rigorous automated and manual validation, before being used to fine-tune backbone LLMs. The performance of Cypher generation is evaluated with a manually annotated evaluation dataset, complemented by a qualitative assessment using GPT as a judge. Additionally, we conduct a scalability test by fine-tuning the LLMs with larger synthetic datasets, which demonstrates that the synthetic data generated using our method does not collapse into simple patterns, thereby establishing the robustness of our approach for larger-scale applications.
SyntheT2C is tested with two medical databases: the LHY database and the Hetionet database (details in Section 5). The generated synthetic dataset, “MedT2C”, will be made public.
In conclusion, our main contributions are:
(1) We propose the SyntheT2C framework containing two pipelines to build synthetic datasets with any Neo4j database. Our method can generate Cypher that are both grammatically correct and syntactically diverse, facilitating the construction of SFT datasets.
(2) We test and validate the effectiveness and scalability of the synthetic dataset generated with SyntheT2C. The LLMs after fine-tuning show improved Cypher writing abilities.
(3) We opensource a synthetic dataset MedT2C of optimal size, ready to be used for SFT.
2 Related works
2.1 Knowledge Graph and graph database
In recent years, KGs have emerged as fundamental resources for organizing, representing, and querying vast amounts of interconnected information or domain-specific knowledge. These graphs find applications across various domains, including but not limited to, healthcare Cui et al. (2024); Abu-Salih et al. (2022), finance Elhammadi et al. (2020); Kertkeidkachorn et al. (2023), and e-commerce Xu et al. (2020). In the realm of Natural Language Processing and Artificial Intelligence (AI), KGs serve as invaluable sources of context and factual knowledge, enabling systems to reason, infer, and generate responses with enhanced accuracy and coherence.
To handle the processing of graph-structured data, a series of graph databases were invented, including Neo4j Neo4j (2012), NebulaGraph Wu et al. (2022), and Amazon Neptune Bebee et al. (2018). Among them, our work focuses on the Neo4j database Neo4j (2012), a widely used graph database management system that excels in modeling and querying highly interconnected data. Neo4j database employs a powerful query language called Cypher for expressing complex graph patterns and retrieving specific data subsets.
2.2 Large Language Models
LLMs are advanced AI models that have been trained on vast amounts of text data to understand and generate human-like language. After the recent breakthrough marked by the release of InstructGPT Ouyang et al. (2022) by OpenAI, a series of LLMs are released, featuring different advantages and drawbacks, e.g., the series of GPT models Brown et al. (2020); OpenAI (2023) by OpenAI, Llama Meta (2024) by Meta, Qwen Bai et al. (2023) by Alibaba Cloud, InternLM Cai et al. (2024b) by Shanghai AI Lab, etc. LLMs can comprehend and generate text across a wide range of topics and writing styles. Recent researches highlight their ability to utilize external existing tools like calculator, search engine, or databases Patil et al. (2023); Nakano et al. (2022); Cai et al. (2024a); Qin et al. (2023). This ability is usually abstracted as “Function calling”, and many of its implementations involve generating codes or queries with LLMs to interact with external tools.
2.3 Code generation
Code Generation is the process of automatically producing executable code from a higher-level representation or natural language. With the advent of LLMs, code generation has experienced a significant advancement. LLMs can now be trained on vast amounts of code and programming-related text materials, enabling them to understand and generate code snippets based on given requirements (e.g., Codex Chen et al. (2021), Polycoder Xu et al. (2022), and Code Llama Rozière et al. (2024)). By leveraging the contextual understanding Dong et al. (2023) and language capabilities of LLMs, code generation becomes more efficient, accurate, and adaptable. Code generation with LLM is not only useful in helping developers to write codes but also in providing a powerful “language” for LLM to interact with other tools: LLMs can be tuned to output executable codes or queries to manipulate external resources. This is the fundamental idea for research in “Function Calling” and Multi-Agent Systems. Current code generation methods rely on two methods for evaluation: either with automatic metrics calculated with an annotated evaluation dataset Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Evtikhiev et al. (2023); Zhou et al. (2023a) or with comparison by a judge (human or powerful LLM like GPT-4) Zheng et al. (2023). Both evaluation methods are used in our work.
3 Methodology
Figure 2: Workflow of two pipelines inside SyntheT2C.
3.1 Preliminaries
The goal of the Text2Cypher task is to automatically translate a query $q$ written in natural language to corresponding Cypher query $c$. With the proposed pipelines $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, a synthetic dataset $\mathcal{S}$ is built to fine-tune the backbone LLM $\mathcal{L}$. The synthetic data is generated and validated with a Neo4j database $\mathcal{B}$ and a series of automatic validators $\mathcal{V}=[\mathcal{V}_{1},\mathcal{V}_{2},...,\mathcal{V}_{5}]$. The synthetic dataset after all the validations is denoted as $\mathcal{S}_{v}$. Using $\mathcal{S}_{v}$, $\mathcal{L}$ is fine-tuned into $\mathcal{L}_{ft}$ The Cypher queries generated by $\mathcal{L}$ (resp. $\mathcal{L}_{ft}$) are noted as $c_{1}$ (resp. $c_{2}$).
3.2 Synthetic dataset generation
Generating the synthetic dataset is not trivial because synthetic data usually has difficulty in balancing grammatical correctness, semantic correctness, node coverage, edge coverage, and Cypher complexity. As a result, we propose a method of generation with two pipelines, as illustrated in Figure 2). The LLM-based prompting pipeline ($\mathcal{P}_{1}$), emphasizes semantic variety, while the template-filling pipeline ($\mathcal{P}_{2}$), focuses on syntactic complexity. By employing these complementary pipelines, we aim to produce a synthetic dataset that captures the nuanced balance of linguistic, semantic, and structural properties.
3.2.1 LLM-based prompting pipeline
This pipeline adopts an idea similar to Knowledge Distillation: to SFT a weaker LLM, we could use the Cyphers generated by stronger LLMs. While off-the-shelf LLMs are typically not optimized for Cypher query generation, as Cypher likely constitutes only a small fraction of their pre-training data, they have nonetheless demonstrated strong in-context learning capabilities Dong et al. (2023). Therefore, half of $\mathcal{S}$ is built by few-shot prompting GPT-4o OpenAI (2023). To further simplify the task of generation and to ensure a higher quality of the generated data, we split the whole generation task into (1) extracting information from the database; (2) determining the question categories; and (3) generating the Cyphers for each category with extracted information.
The workflow for the LLM-based prompting method is delineated in Figure 2 (upper part, $\mathcal{P}_{1}$). Initially, we commence by extracting metadata from the KG stored in the Neo4j database $\mathcal{B}$. This extraction includes sampling example nodes and edges to construct few-shot prompts, along with capturing the schema of the database to facilitate the generation of grounded Cyphers. An illustrative instance of extracted metadata is provided in Appendix A. Subsequently, this metadata serves as a foundational component in all ensuing prompts, ensuring the generation of executable Cyphers. Before initiating the Cypher generation process, a preliminary step involves prompting the LLM to propose potential question categories, thereby mitigating the risk of redundant outputs. The backbone LLM undergoes multiple iterations to propose these question categories, as detailed in the prompt showcased in Appendix B.1. These proposed categories are then consolidated to eliminate duplicates, as instructed in the prompt outlined in Appendix B.3. After the deduplication, GPT-4o is prompted to generate synthetic Question-Cypher pairs with the prompt outlined in Appendix B.2. In our experiment, we fix a list of 12 categories (referred to as categories) to facilitate the comparison.
3.2.2 Template-filling pipeline
The second pipeline of Cypher generation adopts the template-filling method, a classic approach in code generation known for its flexible output and potentially complex syntax. We introduce this pipeline as a complement to the first one, leveraging manually crafted templates to generate Cyphers with more advanced syntax, thereby enabling backbone $\mathcal{L}$ to solve complicated questions.
In this pipeline, depicted in Figure 2 (lower part, $\mathcal{P}_{2}$), numerous templates are initially manually authored. Subsequently, actual values from different fields are sampled from the Neo4j database $\mathcal{B}$ to populate these templates, resulting in the generation of complete executable Cypher queries.
One such template is illustrated in Figure 4. In this example, the subschema is introduced to manage cases where the entire database cannot be loaded at once, necessitating the selection and injection of only the relevant subgraph into the prompt. The variables label_i and prop_j represent the randomly sampled names of nodes and their attributes. These templates are initially crafted taking inspiration from Cypher Generator Onofrei (2024), then enriched and verified by the authors.
Figure 3: Illustration of the automatic validators.
def prompter(label_1, prop_1, prop_2):
subschema = get_subgraph_schema(jschema, [label_1], 2, True)
message = {
"prompt": "Convert the following question into a Cypher query using the provided graph schema!",
"question": f"""Find all {prop_1} for {label_1} that have {prop_2} after January 1, 2020!""",
"schema": f"Graph schema: {subschema}",
"cypher": f"MATCH (n:{label_1}) WHERE date(n.{prop_2}) > date(’2020-01-01’) RETURN n.{prop_1}"
return message
Figure 4: Example template in Template-filling pipeline.
Once these templates are established, synthetic Cyphers with complex syntax can be effortlessly generated. However, it is important to note that crafting and validating these templates require considerable time and effort.
3.3 Quality validation
To ensure the quality of the generated synthetic Question-Cypher pairs before their application in SFT, it’s imperative to conduct thorough validation to prevent the “garbage in, garbage out” scenario. However, manually scrutinizing thousands of Cypher queries is arduous and time-consuming. In response, a suite of automatic validators has been implemented to alleviate the burden of manual inspection. In the end, the Cyphers that pass through these automated validators undergo a final round of meticulous manual validation by researchers.
3.3.1 Automatic validation
We propose five automatic validators: the Grammatical Validator, Semantic Validator, Entity Validator, Schema Validator, and Coherence Validator, each playing a crucial role in ensuring the integrity of the generated synthetic data. These validators’ fundamental concepts are illustrated in Figure 3. The LLM used in the validators is GPT-3.5-Turbo.
The Grammatical Validator validates the syntax correctness of each Cypher in $\mathcal{S}$ by executing them in the deployed graph database $\mathcal{B}$. If a Cypher is executed without encountering any “Error/Exceptions”, it is deemed to have passed this validation.
The design of Semantic Validator is inspired by the research in machine translation Hoang et al. (2018). This validator utilizes an LLM to translate the generated Cypher back into a natural language question. It then computes the semantic similarity between the translated question and the original question. If the similarity score exceeds a predefined threshold, the Cypher passes validation. We also implement an alternative version of the Semantic Validator, where the LLM assesses semantic similarity directly. Both versions produce coherent validation results, with the latter being adopted for efficiency in subsequent experiments. The prompt used in this validator is presented in Appendix C.1.
The Entity Validator assesses the coverage of entities in the generated Cyphers. The entities in the original question $q$ are extracted via Named Entity Recognition (NER) using the spaCy Honnibal and Montani (2017) model en_core_web_sm. Entities in the generated Cypher $c$ are parsed and extracted using Regular Expressions. A successful validation requires 100% coverage of $q$’s entities in $c$. English entities are first transformed into lemmas using spaCy for fuzzy matching.
Subsequently, the Schema Validator ensures the correctness of relations in the generated Cyphers. Relations in $c$ are extracted via Regular Expressions and validated against the schema of $\mathcal{B}$. A Cypher passes this validation only when all contained relations are valid edges.
Lastly, the Coherence Validator executes the Cypher against $\mathcal{B}$ and evaluates the coherence between the execution results and the original question with LLM, using the prompt presented in Appendix C.2.
In the end, only Cyphers that have passed all validations proceed to manual validation.
3.3.2 Manual validation
Each Cypher checked by the validators is randomly assigned to two researchers, who independently assess its quality. If both researchers provide a unanimous judgment, their consensus is adopted. In cases of divergent opinions, a third researcher is brought in for further review. The final validation outcome for such Cyphers is determined through a majority vote among the three researchers.
4 Experiments
4.1 LHY and Hetionet Graph databases
Throughout our experiment, we employed two Neo4j databases housing general medical knowledge in graph form: the LHY Medical Knowledge Database (referred to as “LHY”) and the Hetionet Medical Knowledge Database (referred to as “Hetionet”). Both databases are publicly accessible, differing primarily in language: the data within the LHY database is presented in Chinese, whereas Hetionet is written in English.
The LHY Database Liu (2018) serves as the backend database for a Medical Question-Answering system. This database comprises comprehensive medical knowledge, encompassing a wide array of diseases, symptoms, drugs, and related information. Its content is sourced from medical websites, meticulously cleaned, reorganized, and stored within a Neo4j database. There are about 44k entities and 300k relations in it.
Hetionet Himmelstein et al. (2017) is an open and free-to-use database of biomedical knowledge resource implementing “hetnet” model. Aggregating insights from 29 public databases, Hetionet boasts a knowledge network spanning various fields, encompassing a wide array of entities, including genes, compounds, anatomical structures, diseases, symptoms, side effects, etc. There are approximately 47k entities and 2.2 million relations in the Hetionet database.
The detailed statistics of both databases are presented in Appendix D.
Figure 5: Result of Supervised Fine-Tuning each LLMs with MedT2C. Accuracy annotations marked in white box.
4.2 Evaluation dataset and metrics
We utilize a dataset comprising 300 manually annotated and verified samples to evaluate our experiments. This dataset includes 150 questions annotated based on the Hetionet and LHY databases, respectively. Take Hetionet as an example, for every category among the 12 categories generated in Section 3.2.1, we employ GPT-3.5-Turbo to generate 10 new questions, forming 120 “in-domain” questions. Additionally, we introduce 3 unseen categories and generate 10 new questions for each new category, totaling 30 “out-of-domain” questions. For each of the 300 questions, the authors write a ground-truth Cypher query, which is then executed against the two databases to get the ground-truth execution results.
This annotated dataset allows us to evaluate two aspects of LLMs’ Cypher generating performance:
(1) Cypher quality, which is crucial if the generated Cypher is integrated into larger systems;
(2) Execution result accuracy, to gauge the quality of the output for end users.
4.2.1 Evaluation of Cypher quality
The backbone LLMs, both pre-SFT and post-SFT, are tasked with generating Cyphers for the 300 questions in the evaluation dataset. Using GPT-4o OpenAI (2023), we determine the superior Cypher from the two provided versions. For each pair of Cyphers, we conduct two evaluations by varying the order of presenting the Cypher queries in the prompt to mitigate order-induced bias. If evaluations of both orders yield identical results, this judgment is accepted as the final outcome; otherwise, it is deemed a “Draw”.
4.2.2 Evaluation of execution result accuracy
The generated Cyphers $c_{2}$ are executed on database $\mathcal{B}$ to get execution results $res_{gen}$. Then the accuracy ($acc$) is calculated with the ground-truth execution results $res_{gt}$ like this:
$acc=\frac{\#(res_{gen}\cap res_{gt})}{\#(res_{gen})},$ (1)
where #(.)\#(.)# ( . ) calculates the cardinality of a set.
4.3 Experiment setup
4.3.1 Cypher LLMs
Extensive experiments are conducted with four LLMs, including open/closed-source models. For open-source models, we evaluate Llama3, Qwen2 and InternLM2. For closed-sources model, we test GPT. The exact versions of the backbone LLMs are listed in Appendix E.
4.3.2 Supervised Fine-Tuning
We utilize Low-Rank Adaptation (LoRA) to fine-tune the vanilla LLMs. Specifically, the open-source models are trained for 6 epochs with a linear scheduler, starting at a learning rate of 1e-6. AdamW is used as the optimizer, and the training batch size is 6. The fine-tuning of GPT is facilitated by its official API. The experiments on all LLMs, are conducted on Nvidia GeForce 4090 GPU. All the experiments totaled about 1100 GPU hours.
4.4 Supervised Fine-Tuning experiments
The backbone LLMs are fine-tuned with the MedT2C dataset, comprising 750 samples generated with the two pipelines and two Neo4j databases, totaling 3000 samples. The MedT2C dataset contains high-quality Question-Cypher pairs that passed all the automatic validations as well as the manual validation. In Appendix F we report the passing rate of each validator as a guide for further improvement of MedT2C’s data quality.
A list of LLMs including GPT, Llama, Qwen, and InternLM are fine-tuned using MedT2C. We evaluated the change in Cypher writing performance of these LLMs, and the results are shown in Figure 5. The results show that MedT2C helps the LLM to produce more Cypher queries that are on par with or better than the human annotated ones.
In Figure 5, the win rates are calculated in comparison with the ground-truth Cyphers. We further conduct an experiment to compare directly the $c_{1}$ and $c_{2}$ generated with the same LLM with GPT-4o, using the prompt presented in Appendix G. The comparison results are shown in Figure 6. From these results, we can conclude that while the improvement may appear minor when comparing against the ground-truth Cyphers, a visible enhancement in Cypher quality is evident when comparing to the Cyphers generated by the pre-SFT model. We explain this difference as follows: the human annotations have a far higher quality than the Cyphers generated by vanilla LLMs. Therefore, even though the LLMs are enhanced after SFT, their output is still inferior to the human-annotated Cyphers, which is why the evaluation results in Figure 5 seem largely unchanged.
Figure 6: Impact of SFT on each LLM. The Cypher generated with pre-SFT and post-SFT LLMs are compared directly with GPT-4o.
4.5 Scaling experiments
In this section, we test the scalability of our pipeline for generating synthetic data. We rerun the data generation pipelines to create scaled versions that are 1/16, 1/4, 4, 8, and 16 times the size of the original MedT2C. Vanilla LLMs are then fine-tuned with these scaled datasets. The results are reported in Figure 7. These results demonstrate that, up to the size of the MedT2C dataset, increasing the size of the synthetic dataset leads to better performance, especially in terms of Cypher Quality. However, once the size exceeds that of MedT2C, further increasing the dataset size results in either marginal improvements or decreases. Based on this experiment, we determine the optimal size for the published MedT2C dataset (highlighted in red), as it balances efficiency and performance.
Figure 7: Plots of scaling test’s results.
4.6 Ablation experiments
To evaluate the efficacy of each component introduced, we conduct a series of ablation experiments. First, we test the pipelines by running SFT experiments using only the data generated by each pipeline individually. We then verify the effectiveness of each automatic validator by evaluating them in isolation, using only one validator at a time. Since each component is designed to be modular and independent, we adopt this mode of ablation, rather than removing the components one by one from the complete setting, to emphasize the increment brought by each component separately. For both ablation tests, the backbone LLM is fixed as Llama3. The dataset size is set to be the same as MedT2C (3000 in total). The experiments results are reported in Table 1 and Table 2 respectively. Here the Cypher Quality is calculated with respect to ground-truth Cyphers.
Settings Cypher Quality Result Acc.
Pre-SFT 38.67%(–) 27.83%(–)
   LHY-LLM 41.67%(+3.00) 27.86%(+0.03)
   LHY-Temp. 34.67%(-4.00) 26.54%(-1.29)
   Hetionet-LLM 42.83%(+4.16) 33.09%(+5.26)
   Hetionet-Temp. 36.00%(-2.67) 26.68%(-1.15)
All (MedT2C) 44.00%(+5.33) 39.65%(+11.82)
Table 1: Results of pipeline ablation test.
As presented in Table 1, when we use only the data generated by the template-filling pipeline to SFT the Llama3 model, the model’s performance actually declined after SFT. This can be attributed to the design of the template-filling pipeline, which emphasizes generating syntactically complex Cypher queries. When SFT is performed using only this data, the backbone LLM tends to produce unnecessarily complicated Cypher queries (e.g., breaking one query into two and then merging them). While this ability to write more complex Cypher queries is not directly reflected in the evaluation metrics, as the “hard questions” requiring advanced syntactic knowledge constitute only a small portion of the evaluation dataset, such data can enhance the LLM’s generalization capacity when combined with data from the LLM-based prompting pipeline.
Settings Cypher Quality Result Acc.
Pre-SFT 38.67%(–) 27.83%(–)
   No validator 38.34%(-0.33) 27.96%(+0.13)
   ✓Grammar V. 38.34%(-0.33) 28.95%(+1.12)
   ✓Semantic V. 43.67%(+5.00) 31.65%(+3.82)
   ✓Entity V. 40.00%(+1.33) 28.03%(+0.20)
   ✓Schema V. 42.00%(+3.33) 26.11%(-1.72)
   ✓Coherence V. 41.33%(+2.66) 32.05%(+4.22)
All (MedT2C) 44.00%(+5.33) 39.65%(+11.82)
Table 2: Results of validator ablation test.
As shown in Table 2, each individual validator contributes some improvement, either in terms of Cypher quality or the accuracy of the execution results. Notably, the combination of all five validators yields the most significant increase in performance. This can be attributed to the validators’ collective ability to mitigate the majority of the bugs in the SFT dataset, thereby enhancing the overall quality of the generated Cypher queries.
5 Limitations
The primary limitation of our work is the challenge in writing the templates. While the templates are designed to be independent from the base Neo4j databases, some adaptation work is still necessary when applying them to new Neo4j databases. Additionally, writing new templates is time-consuming, making the expansion of the current template library difficult. Furthermore, a significant number of the generated Cypher queries are directly filtered out during the construction of MedT2C, resulting in a waste of resources. Developing methods to quickly fix Cyphers that do not satisfy all the validation criteria, instead of simply regenerating more Cyphers, could help reduce the carbon footprint of SyntheT2C.
6 Potential risks
Even though SyntheT2C is designed to automatically generate synthetic datasets, its usage requires close monitoring and manual validation to prevent the inadvertent inclusion of private or sensitive information. Additionally, the post-SFT LLM should be used with caution. Despite improvements in their Cypher generation performance, there remains a slight risk of producing embedded Cyphers that could lead to issues such as Out-of-Memory.
7 Conclusion
We present SyntheT2C, a comprehensive framework to generate synthetic data for SFT various LLMs on the Text2Cypher task. Our approach encompasses dataset construction, data validation, and SFT evaluation, providing a reference framework for future research in the Cypher-related field. Additionally, our findings confirm the effectiveness of synthetic data, suggesting that similar techniques can address problems where annotation is difficult or insufficient. Finally, we will also open-source the MedT2C dataset, aiming to contribute to the technical advancements in relevant topics.
Appendix A Example of extracted KG information
Here we present the information (metadata) extracted from the KG database “Hetionet” in Figure 8. We stored the metadata of the KG, including the node properties, the relationship properties, and the valid relationships. This information is integrated in the following prompts to ensure that the LLM output is correct Cyphers. In other prompts, this metadata is referred to as schema.
Appendix B Prompts for LLM-based prompting pipeline
In this appendix, we present all the prompts we used in the LLM-based prompting pipeline.
B.1 Prompt to propose categories of questions
In Figure 9 we show the prompt used to propose candidate categories of questions. We decided to first generate categories of questions instead of generating directly the questions because this practice helps reduce duplicated questions.
B.2 Prompt to generate questions for each category
This prompt presented in Figure 11 is used to generate questions in natural language for each proposed category. This prompt includes few-shot examples to help ensure the output Cypher follows the format requirements.
B.3 Prompt to merge categories of questions
The prompt presented in Figure 10 is used to merge the previously generated categories. The merged and de-duplicated list of categories is then stored and will be referred to as category in later prompts.
Appendix C Prompts used in automatic validators
C.1 Prompt of Semantic Validator
Here we present the prompt used in the Semantic Validator in Figure 12. The schema mentioned in this prompt is the metadata presented in Appendix 8. The example represents the few-shot examples written by the authors, here we show the English example for the Hetionet database in Figure 13. Lastly, the json_object in the prompt contains the question and the Cypher query to be evaluated.
C.2 Prompt of Coherence Validator
In this appendix, we present the prompt used in the Coherence Validator in Figure 14. Similar to other prompts, we provided few-shot examples in this prompt. The question and results in the prompt are the original question and execution results used as the input for this validation.
Appendix D Important statistics of the LHY and the Hetionet databases
Here we present the important statistics of the LHY database in Table 3 and Table 4, including the examples of nodes and entities inside this database. The examples in both tables are translated from Chinese to English. Similarly, the important statistics of the Hetionet database with examples of nodes and entities are grouped in Table 5 and Table 4.
Appendix E Exact versions of the backbone LLMs
The exact versions of the LLMs used in our experiments are listed in Table 7. Except GPT-3.5-Turbo, the backbone LLMs are deployed locally using the versions available on HuggingFace.
Appendix F Passing rate of MedT2C for each automatic validator
The passing rate of MedT2C dataset for each automatic validator is reported in Table 8. The LLM used in the Semantic Validator and the Coherence Validator is GPT-3.5-Turbo. These two validators are not run on the LLM-based prompting pipeline because this pipeline uses GPT-4o. Given that GPT-4o is more powerful than GPT-3.5-Turbo, it is not accurate to evaluate its output with GPT-3.5-Turbo, nor with GPT-4o itself. Besides, noted that the passing rate of Coherence Validator is especially low compared to other passing rate. This is because for Coherence Validator specifically, the samples that failed any one of the previous validators is judged as False directly to save the calling of GPT API. Therefore the passing rate of Coherence Validator reported here is lower than the actual one, but it does not affect the “All passed” ratio.
Appendix G Prompts used for Cypher quality evaluation
We use GPT-4o to judge the quality of two versions of Cypher queries corresponding to the same set of questions written in natural language. The prompt used for this part is shown in Figure 15. We provide different aspects of evaluation and ask GPT-4o to give detailed reasons when evaluating, because these techniques bring more accurate evaluation results in practice.
Node properties are the following:
Disease {easy_get: STRING, cure_lasttime: STRING, cured_prob: STRING, name: STRING, desc: STRING, prevent: STRING, cure_way: LIST, cause: STRING, cure_department: LIST},Drug {name: STRING},Food {name: STRING},Check {name: STRING},Department {name: STRING},Producer {name: STRING},Symptom {name: STRING}
Relationship properties are the following:
recommand_eat {name: STRING}, no_eat {name: STRING}, do_eat {name: STRING}, belongs_to {name: STRING}, common_drug {name: STRING}, drugs_of {name: STRING}, recommand_drug {name: STRING}, need_check {name: STRING}, has_symptom {name: STRING}, acompany_with {name: STRING}
The relationships are the following:
(:Disease)-[:belongs_to]->(:Department), (:Disease)-[:common_drug]->(:Drug), (:Disease)-[:recommand_drug]->(:Drug), (:Disease)-[:need_check]->(:Check), (:Disease)-[:has_symptom]->(:Symptom), (:Disease)-[:acompany_with]->(:Disease), (:Disease)-[:recommand_eat]->(:Food), (:Disease)-[:no_eat]->(:Food), (:Disease)-[:do_eat]->(:Food), (:Department)-[:belongs_to]->(:Department), (:Producer)-[:drugs_of]->(:Drug)
Figure 8: The metadata extracted from the Hetionet database.
You are an experienced and useful Python and Neo4j/Cypher developer.
I have a knowledge graph for which I would like to generate interesting questions that span 12 categories (or types) about the graph. They should cover single-node questions, two or three more nodes, relationships, and paths. Please suggest 12 categories together with their short descriptions. Here is the graph schema:
{schema}
Figure 9: The prompt used to generate categories of questions.
You are an experienced doctor and you have a knowledge graph for which you would like to generate interesting questions of 12 categories.
Here are some candidate categories:
{categories_list}.
You should merge similar categories and remove the duplicates. Finally, give me a short description of each category.
Figure 10: The prompt used to merge proposed categories.
You are an experienced Cypher developer and English-speaking doctor and a helpful assistant designed to output JSON
Generate {k} questions and their corresponding Cypher statements about the Neo4j graph database with the following schema:
{schema}
The questions should cover {category} and should be phrased in a natural conversational manner. Make the questions diverse and interesting. Make sure to use the latest Cypher version and that all the queries are working Cypher queries for the provided graph. You may add values for the node attributes as needed. Do not add any comments, do not label or number the questions.
Here are some examples of the Question-Cypher pairs to be generated:
"question": "What are the diseases that commonly accompany Depression’?",
"cypher": "MATCH (d1:Disease {{name:’Depression’}}) -[:acompany_with]-> (d2:Disease) RETURN d2.name"
"question": "Can you list diseases that commonly accompany Cancer’?",
"cypher": "MATCH (d1:Disease {{name:’Cancer’}}) -[:acompany_with]-> (d2:Disease) RETURN d2.name",
Now its your turn to generate the question and Cypher pairs:
Figure 11: The prompt used to generate questions.
Ent. Type # Ent. Examples
Check 3,353 Bronchography
Department 54 Department of Plastic and Reconstructive Surgery
Disease 8,807 Thrombosed Vasculitis
Drug 3,828 Jingwanhong Hemorrhoid Cream
Food 4,870 Tomato and Vegetable Beef Ball Soup
Producer 17,201 Tongyi Pharmaceutical Penicillin V Potassium Tablets
Symptom 5,998 Hypertrophy of breast tissue
Total 44,111 /
Table 3: Entities in LHY Database.
Rel. Type # Rel. Examples
belongs_to 8,844 <Gynaecology, belongs_to, Obstetrics and Gynaecology>
common _drug 14,649 <Yang Qiang, common_drug, Phentolamine mesylate dispersible tablets>
do_eat 22,238 <Thoracic spine fracture, do_eat, Blackfish>
drugs_of 17,315 <Penicillin V Potassium Tablets, drugs_of, Tongyi Pharmaceuticals Penicillin V potassium tablets>
need _check 39,422 < Unilateral emphysema, need_check, Bronchography>
no_eat 22,247 <Lip disease, no_eat, Almonds>
recommend_drug 59,467 <Mixed hemorrhoids, recommend_drug, Jingwanhong Hemorrhoid Cream>
recommend_eat 40,221 <Synovial effusion, recommend_eat, Beef Ball Soup with Tomato and Vegetable Punch>
has_ symptom 5,998 <Early Breast Cancer, has_symptom, Hypertrophy of breast tissue>
accompany _with 12,029 <Valvular insufficiency of the traffic veins of the lower extremities, accompany_with, Thromboembolic vasculitis>
Total 294,149 /
Table 4: Relations in LHY Database.
Ent. Type # Ent. Examples
Anatomy 402 Digestive System
Biological_process 11,381 Protein Sialylation
Cellular_component 1,391 Meiotic Spindle
Compound 1,552 Mannitol
Disease 137 Hypertension
Gene 20,945 STRIP2
Molecular_function 2,884 Vitamin Transporter Activity
Pathway 1,822 Glycolysis
Pharmacologic_class 345 Decreased Blood Pressure
Side_effect 5,734 Subileus
Symptom 438 Ageusia
Total 47,031 /
Table 5: Entities in Hetionet Database.
Rel. Type # Rel. Examples
Anatomy–downregulates–Gene 102,240 <Bronchus, downregulates, GRIA2>
Anatomy–expresses–Gene 526,407 <Myocardium, expresses, EFHD1>
Anatomy–upregulates–Gene 97,848 <Adipose tissue, upregulates, PARM1>
Compound–binds–Gene 11,571 <Sildenafil, binds, CYP3A4>
Compound–causes–Side_Effect 138,944 <Ciprofloxacin, causes, Visual Disturbance>
Compound–downregulates–Gene 21,102 <Tacrolimus, downregulates, UBE2C>
Compound–palliates–Disease 390 <Fluvoxamine, palliates, Panic Disorder>
Compound–resembles–Compound 6,486 <Clotrimazole, resembles, Bifonazole>
Compound–treats–Disease 755 <Reserpine, treats, Hypertension>
Compound–upregulates–Gene 18,756 <Estriol, upregulates, KLHL9>
Disease–associates–Gene 12,623 <Parkinson’s Disease, associates, HTR7>
Disease–downregulates–Gene 7,623 <Schizophrenia, downregulates, MLST8>
Disease–localizes–Anatomy 3,602 <Migraine, localizes, Brain>
Disease–presents–Symptom 3,357 <Lung Cancer, presents, Constipation>
Disease–resembles–Disease 543 <Bone Cancer, resembles, Head and Neck Cancer>
Disease–upregulates–Gene 7,731 <Malaria, upregulates, JAK2>
Gene–covaries–Gene 61,690 <IMP3, covaries, OR8U8>
Gene–interacts–Gene 147,164 <TRIM27, interacts, MED21>
Gene–participates–Biological_Process 559,504 <ABCA1, participates, Lipid Homeostasis>
Gene–participates–Cellular_Component 73,566 <KLHL14, participates, Neuronal Cell Body>
Gene–participates–Molecular_Function 97,222 <TOP2B, participates, ATPase Activity>
Gene–participates–Pathway 84,372 <GGT5, participates, Metabolism>
Gene-regulates-Gene 265,672 <BCCIP, regulates, HLTF>
Pharmacologic_Class–includes–Compound 1,029 <Allergens, includes, Benzocaine>
Total 2,250,197 /
Table 6: Relations in Hetionet Database.
You are an experienced Cypher developer, English Master, and a helpful assistant that helps me to verify whether the cypher is coherent with the question!
You will be given a JSON object containing a question and a cypher query. You should first take a look at the schema provided below. The schema is the graph database on which the cypher queries will be run.
The schema:
{schema}
You must organize your answer step by step and in the end, you should make your judgment.
Here are three areas that you should pay attention to:
1. whether the output of cypher is coherent with the question, which means that the output of cypher must contain the information that the question asks.
2. If the question points out a piece of key information, you should check whether this key information is pointed out in the cypher. For example, if the question provides a piece of exact information such as the exact name of the disease, this information can not be inconsistent in the cypher. If there is no exact key information, you can skip this area.
3. whether this cypher answers the question provided in the JSON object. You should simulate the cypher step by step according to the schema provided. Then you should judge whether this cypher is in line with the question.
You should make your judgment according to these three areas. If there are no problems in these three areas in the cypher, you must answer with True’. Otherwise, you should answer with False’.
Here are some example JSON objects:
{example}
Now its your turn to answer! Here is the JSON object you should evaluate:
{json_object}
Now evaluate carefully the JSON object and provide your answer step by step.
Figure 12: The prompt used in Semantic validator.
<|Example 1|>
"question": "Which diseases belong to the Psychiatry department?",
"cypher": "MATCH (d:Disease)-[:belongs_to]->(dept:Department) WHERE dept.name = Neurology RETURN d.name"
<|Answer 1|>
The cypher is not in line with the question because the question is to find the diseases in the Psychiatry department but the department name in the cypher is Neurology department.
Since the key information is inconsistent, I would mark this JSON object as False.
<|Example 2|>
"question": "Which foods should be avoided for the disease Brain tumor’?",
"cypher": "MATCH (d1:Disease {name:’Brain tumor’})-[:no_eat]->(d2:Food) RETURN d2.name"
<|Answer 2|>
Firstly, the output of cypher contains the key information the food asked by the question.
Secondly, the key information Brain tumor provided in the question is contained in the cypher.
Finally, the logic of cypher is exactly similar to the question.
So, I think this JSON object is True.
<|Example 3|>
"question": "What pathways do the genes BRCA1 and BRCA2 participate in?",
"cypher": "MATCH (g:Gene)-[:PARTICIPATES_GpPW]->(:Pathway) WHERE g.name IN [’BRCA1’, BRCA2’] RETURN g.name"
<|Answer 3|>
There are two errors.
Firstly, as the question asks for pathways but the output of cypher is the name of the gene, the output of the cypher is inconsistent with the question.
Secondly, the question is to find the pathway that both the genes BRCA1 and BRCA2 participate in. But the cypher matches the pathways that BRCA1 or BRCA2 participates in. The logic AND and OR are totally different.
Therefore, I think this JSON object is False.
Figure 13: The English few-shot examples used in the Semantic Validator.
You are an experienced medical assistant who has mastered English and medical knowledge.
You will be given a question and the responses given by the doctor. The doctor is very professional, he gives direct responses. But he sometimes misunderstands the problem. Your task is to check if the results are coherent with the question by analyzing the category.
For example, if the question asks for food and the answer is food, in this case, it is relevant because the category is the same. Even if the foods dont seem to be directly related, you can not deny them because the doctor is professional.
But if the question asks for food, the doctor gives the response on sports. You should point out this error because the category is different.
As a medical assistant, you just need to pay attention to whether the category of the answer corresponds to the category that the question asks. You dont need to think about the reasonableness of the answer.
Answer with True if the category is the same. Otherwise, answer with False’.
You need to carefully explain your answer.
Here are some examples of questions and results:
<Example 1>
Question: Find out the diseases associated with the Oncology department.
Responses by the doctor: Breast cancer, Pancreatic cancer, Colon cancer
Your reply: Breast cancer, pancreatic cancer, and colon cancer belong to the Oncology department. And the question asks for diseases. So I think it is relevant, and my answer is True.
<Example 2>
Question: Which foods should be avoided for the disease Coeliac disease’?
Responses by the doctor: Swimming, Running, Biking, Walking
Your reply: The responses are sports. But this question asks for food. So I think it is not relevant, my answer is False.
Now its your turn to verify if the responses are relevant to the question.
Remember! You just need to pay attention to whether the answer corresponds to the question. You dont need to think about the reasonableness of the answer.
Question:{question}
Responses by the doctor: {results}
Your reply:
Figure 14: The prompt used in Coherence Validator.
LLM name LLM version LLM site
GPT gpt-3.5-turbo-16k https://platform.openai.com/docs/models/gpt-3.5-turbo
Llama3 Meta-Llama-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
InternLM2 internlm2-7B https://huggingface.co/internlm/internlm2-base-7b
Qwen2 Qwen2-7B https://huggingface.co/Qwen/Qwen2-7B
Table 7: Versions of the backbone LLMs
Database Pipeline Grammatical Validator Semantic Validator Entity Validator Schema Validator Coherence Validator All passed
LHY LLM-based prompting 99.69% N/A 99.62% 82.77% N/A 83.87%
LHY Template-filling 99.87% 92.34% 100% 99.87% 28.59% 27.21%
Hetionet LLM-based prompting 96.08% N/A 99.08% 61.69% N/A 64.79%
Hetionet Template-filling 100% 91.81% 99.52% 100% 38.15% 36.66%
Table 8: MedT2C’s passing rates of each automatic validator.
You are an expert in medical field and Cypher query language. You are asked to evaluate the quality of the Cypher queries generated by 2 models for the same question. You will be first given the question written in natural language. Then you will be given the Cypher queries generated by 2 models. Your task is to compare the quality of these two Cyphers and select the better one. You should consider the following aspects when selecting the better Cypher:
1. Syntactical correctness: whether the Cypher query is syntactically correct;
2. Semantic correctness: whether the Cypher query can correctly answer the question;
3. Readability: whether the Cypher query is easy to read and understand;
4. Efficiency: whether the Cypher query is efficient in terms of time and space complexity;
5. Conciseness: whether the Cypher query is concise and clear;
6. Completeness: whether the Cypher query can cover all the necessary information in the database.
You should select the better Cypher query based on these aspects. Output your selected Cypher as well as your reasons.
Here is the question:
"question": "{{ question }}"
Here are the outputs of the models:
"number": "1",
"cypher": "{{ cypher_1 }}"
"number": "2",
"cypher": "{{ cypher_2 }}"
Your output should be in the following format, DO NOT output anything other than this JSON object:
"better_cypher": "1",
"reason": "reasons why 1 is selected"
Now select the better Cypher and give your reasons:
Figure 15: The prompt used in Cypher quality evaluation.